The legs of a right triangle are 3 and 8 centimeters long, respectively. What is the length of the hypotenuse? If necessary, round your answer to two decimal places.

Answer by checkley77(12844)

3^3+8^2=c^2
9+64=c^2
73=c^2
c=sqrt73
c=8.544 ans.